OpenStudy (anonymous):

a function "f" is defined as follows: f:[-2,a] > R, where f(x) = (x-1)squared - 4. if f(a) = 12, find the value of a??

OpenStudy (anonymous):

sorryi meant if f(x) = 12, what is x.

OpenStudy (anonymous):

no wait the first one was right. i got things mixed up ><

OpenStudy (anonymous):

This looks worse than it is. You're given the function f(x) and you know that when x=a, the function equals 12; that is, for x=a, f(a)=12. So,\[f(a)=(a-1)^2-4=12\]Solving for a:\[(a-1)^2=16\rightarrow a-1=\pm4 \rightarrow a=-3, 5\]So a can be either -3 or 5.

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

ur welcome. feel free to 'fan' me.

OpenStudy (anonymous):

oh okay. i just did. im new to this hahaha,

OpenStudy (anonymous):

cheers :)